# Multi-Order Hyperbolic Graph Convolution and Aggregated Attention for Social Event Detection

### v1

@@ 5.3.2 Parameter Settings
All the experiments with the MOHGCAA model were conducted on an NVIDIA V100 GPU with 12 24-core Intel Xeon processors.
@@ 5.3.4 Performance Comparison of the MOHGCAA Model (Answer Q1)
euclidean Word2vec 0.34 0.44 0.13 0.02
LDA 0.22 0.29 0.04 0.01
WMD - 0.65* 0.5* 0.06*
BERT 0.53 0.64 0.44 0.07
KPGNN - 0.7* 0.52* 0.22*
FinEvent - 0.8* 0.69* 0.48*
hyperbolic HNN 0.89 0.92 0.73 0.88
@@ 5.4.2 Performance Comparison Among Hyperbolic Space Models (Answer Q4)
6 Concussion
In this paper, we addressed the challenges of social event detection (SED) by proposing a novel framework, Multi-Order Hyperbolic Graph Convolution with Aggregated Attention (MOHGCAA). Our approach effectively captures the hierarchical and higher-order relationships in event detection data, overcoming the limitations of existing methods in Euclidean and hyperbolic spaces. Experimental results on multiple datasets demonstrated that MOHGCAA consistently achieves superior performance under both supervised and unsupervised settings, validating its effectiveness and robustness. These findings highlight the framework’s capability to enhance SED applications across diverse scenarios, offering valuable insights for businesses and public services alike. Moving forward, our work opens new possibilities for leveraging hyperbolic graph convolutional networks in other hierarchical data domains. Future research may focus on further optimizing the framework for scalability and exploring its potential in real-time event detection and other dynamic graph-based tasks.

### v2

@@ 5.3.2 Parameter Settings
All the experiments with the model were conducted on an NVIDIA GeForce RTX 3090 GPU with AMD EPYC 7302 64-core CPU processors. For other experimental hyperparameters, please refer to Table 5.
@@ 5.3.4 Performance Comparison of the MOHGCAA Model (Answer Q1)
euclidean Word2vec 0.32 0.41 0.12 0.02
LDA 0.2 0.28 0.04 0.01
WMD - 0.63* 0.49* 0.06*
BERT 0.51 0.61 0.41 0.07
KPGNN - 0.69* 0.5* 0.21*
FinEvent - 0.79* 0.69* 0.48*
hyperbolic HNN 0.89 0.91 0.71 0.87
@@ 5.4.2 Performance Comparison Among Hyperbolic Space Models (Answer Q4)
6 Conclusion
In this paper, we addressed the challenges of social event detection (SED) by proposing a novel framework, Multi-Order Hyperbolic Graph Convolution with Aggregated Attention (MOHGCAA). Our approach effectively captures the hierarchical and higher-order relationships in event detection data, overcoming the limitations of existing methods in Euclidean and hyperbolic spaces. Experimental results on multiple datasets demonstrated that MOHGCAA consistently achieves superior performance under both supervised and unsupervised settings, validating its effectiveness and robustness. These findings highlight the framework’s capability to enhance SED applications across diverse scenarios, offering valuable insights for businesses and public services alike. Moving forward, our work opens new possibilities for leveraging hyperbolic graph convolution networks in other hierarchical data domains. Future research may focus on further optimizing the framework for scalability and exploring its potential in real-time event detection and other dynamic graph-based tasks.
Appendix A My Appendix
[…]